OpenStudy (anonymous):

is 5e^(5x) * e^(-x) = 5e^(4x)?

OpenStudy (lgbasallote):

\[\large 5e^{5x} \times e^{-x} \implies 5e^{5x + -x} \implies 5e^{5x - x} \implies 5e^{4x}\] does that answer your question?

OpenStudy (unklerhaukus):

\[a^n\times a^b=a^{a+b}\]\[ 5\times e^{5x} \times e^{-x}=5e^{5x-x}\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

one last thing...

OpenStudy (anonymous):

is 6e^0 = 6?

OpenStudy (unklerhaukus):

\[a^0=1\]

OpenStudy (lgbasallote):

e^0 =1 6(1) = 6

OpenStudy (anonymous):

thank you

OpenStudy (unklerhaukus):

\[\checkmark\checkmark\checkmark\]